### v1

@@ 4.1 Source extraction strategy
The continuum images contain multiscale structures ranging from beam scales (02–03) to maximum recoverable scales (7″–14″), and exhibit a high intensity dynamic range up to 1000 (c.f. Col. 5 in Table 2). The getsf (v231026) extraction algorithm (Men’shchikov 2021) was adopted to identify and characterize sources, as it can identify sources from multi-scale and also from multi-band data. We adopted the getsf definition of sources, which are roundish and significantly stronger than local surrounding fluctuations (of background and noise). The getsf algorithm is recommended for interferometric data, as it effectively handles various backgrounds from filamentary structures to negative bowls originated from imperfect uv-sampling and can separate blended sources in crowded fields like protoclusters (Xu et al. 2023; Cheng et al. 2024).
Our getsf procedure was designed to take full advantages of the dual-band observations. Details of the settings and the main workflow can be found in Appendix B. Here we only provide a brief overview. The inputs are the 1.3 mm and 3 mm continuum images at the original resolutions to avoid potential loss of detection due to the smoothing process. The images are without primary beam corrections for a more uniform noise distribution. The extraction mask includes regions where the primary beam response is greater than 0.3 (pb0.3), to exclude ambiguous detections on the edges. The minimum source size is set as the original resolution, and the maximum source size was set as five times the 1.3 mm beam size, 1.2 (approximately 0.05 pc or 10,000 au at the CMZ distance). The choice of the maximum source size can not only effectively exclude those extended sources that are not core-like structures, but also preserve the completeness of the core-like source catalog. The main getsf procedure was divided into three rounds: preparing flattened components across the two bands; performing single-band extractions for monochromatic catalogs; rerunning a dual-band detection to obtain a combined catalog.
@@ 4.2 Catalog description
Our procedure (see Sect. 4.1) produced two types of source catalogs: mono-band and dual-band. The mono-band source catalogs (mbcat) provide detections and measurements at individual bands, while the dual-band source catalogs (dbcat) combine sources 333getsf accumulates the clean images (background removed) over dual bands by a natural weight as described by Eq. (29) in Men’shchikov (2021). decomposed from the two bands for detection and measurement. In the following, we will introduce the mono-band catalogs in Sect. 4.2.1 and the dual-band catalogs in Sect.4.2.2, and discuss their caveats and complementarity in Sect. 4.2.3. The source catalogs are publicly available at 10.5281/zenodo.15063109.
@@ 4.2.1 Mono-band source catalogs
The 1.3 mm and 3 mm mbcats are presented in Tables 4 and 5, respectively. The significance value (Sig) is equivalent to the signal-to-noise ratio (SNR) of a source on the scale where it is best detected by getsf. The sources are sorted by the Sig value in the descending order, as indicated by the ID in the table. Throughout the DUET series, we use the nomenclature of “{cloud name}-{band}{#ID}”, where “cloud name” refers to either of cloude, SgrC, and 20kms. For example, SgrC-1mm#1 refers to the source with ID of 1 in the 1.3 mm mbcat in the Sgr C cloud, and 20kms-3mm#2 refers to the source with ID of 2 in the 3 mm mbcat in the 20 km s-1 cloud. Following the convention from Motte et al. (2018b), Sig greater than 7 corresponds to a robust detection. There are 437 and 278 robust detections at 1.3 mm and 3 mm in mbcats, corresponding to about 77% and 84% of the monochromatic detections, respectively. To obtain the intrinsic source sizes, we deconvolved the synthesized beams from the observed sizes by the analytical solution introduced in Appendix C. The deconvolved FWHM sizes were converted into physical scales at the distance of 8.277 kpc. The peak intensity and total flux were both primary beam corrected; the total flux of source is aperture corrected by the radio beam shape rather than directly summarizing flux per pixel that belongs to the source.
We do not perform any further filtering criteria based on their intensity profiles, to maintain mbcat’s completeness for future studies at individual bands. For example, some diffuse and extended sources can be analogous to nearby prestellar cores with sizes of 10,000 au (see Table 4 of Xu et al. 2024a, and references therein). The intensity profile of these prestellar cores should be much shallower than typically observed in protostellar cores because of a shallower density profile and a positive temperature gradient (colder in the inner part). Due to the missing flux effect, prestellar cores may appear even fainter than expected.
@@ 4.2.2 Dual-band source catalogs
The dbcat catalogs were obtained in the third round of getsf procedure (see Appendix B), where sources were identified and measured with the combination of both 1.3 mm and 3 mm continuum emissions. Since the fields of view (FoVs) of the two bands only partially overlap, the initial dbcats include sources outside the pb=0.3 mask of one of the bands. We removed such sources to obtain a cleanest source sample within the overlapped FoVs. Finally, a total of 450 sources (including 62 in cloud e, 182 in Sgr C, and 206 in the 20 km s-1 cloud) were included in the dbcat, which are listed in Table 6. The complete version of dbcat is shared in 10.5281/zenodo.15068333. Throughout the paper, we use the nomenclature of “{cloud name}-{db}#ID”. For example, SgrC-db#1 refers to the source with ID of 1 in the dbcat in the Sgr C cloud. As an example, the spatial distribution of the identified sources in the Sgr C cloud is shown in Fig. 3. The complete version for all the three clouds are shared both in an interactive mode (https://xfengwei.github.io/magnifier/index.html) and in 10.5281/zenodo.15068333.
The dbcat sources have undergone a similar filtering criteria as the mbcat ones, but the goodness and significance are combined from both 1.3 mm and 3 mm. Therefore, some dbcat sources could have a small Sig value in one band because of a significant detection at the other band. For example, SgrC-db#30 is marginally detected at 1.3 mm but filtered by Sig and goodness criteria, while it has SNR65 at 3 mm. Another example is SgrC-db#106, which has a SNR17 detection at 1.3 mm but marginal detection at 3 mm. In the complete version of Table 6, the monochromatic Sig values at 1.3 mm and 3 mm are also given.
The dbcat can also resolve crowded sources thanks to the higher angular resolution at 1.3 mm. We refer to the cloud e-Main as an example. We identified cloude-1mm#1, 15, 21, 23, 28, 29, 55, and 61 from a clustered region at 1.3 mm, but only cloude-3mm#1, 16, and 22 at 3 mm due to its coarser resolution. However, all the eight sources were successfully separated and identified in the dbcat.
Source sizes and fluxes at each band were measured within individual monochromatic footprints, so the sizes can differ between the two bands. Two data columns for 1.3 mm and 3 mm respectively were include in the catalog, each of which contains the same parameters as the mbcats. The intrinsic source sizes of both bands were obtained by deconvolving the synthesized beams from the observed sizes (see Appendix C). In Table 6 listed are the major and minor FWHM sizes, position angle, intrinsic source size, peak intensity, and integrated flux over the source.
@@ 4.2.3 Completeness and robustness of the catalogs
As shown by orange and yellow solid lines in Fig. 1, the FoVs at 1.3 mm and 3 mm only partially overlap. The mbcats contain all the detections within the FoVs at the individual bands, while the dbcat only includes sources appearing in the overlapped FoVs of the two bands. In this sense, the mbcat ensures the completeness of the sources within the FoV at each individual band.
Meanwhile, dbcat considers both bands simultaneously for detection, thus raising the robustness of source detection. For example, the mass sensitivity for optically thin dust at 20 K at 1.3 mm is about 4 times better than 3 mm, so a faint dusty source can have robust detection at 1.3 mm but less than at 3 mm. In this case, the 1.3 mm mbcat will include this source, but the dbcat will not. Therefore, the 1.3 mm mbcat catalog is more complete than the dbcat at the low-flux end. In turn, sources in the dbcat are more robust, because they are verified in both bands. This is particularly important for interferometric images, where there can be sidelobes or artifacts from cleaning. Compared to the mbcat, the dbcat combines images to filter out false detections, resulting in a smaller but more robust source catalog.
However, it should be noted that if a source is marginally detected and filtered out at one band but has robust detection at the other, the dbcat will contain it. It usually happens in two cases. The first case is also for optically thin dust core as mentioned above, but the difference happens when 1.3 mm detection reaches as much as 20 but 3 mm is still below 5. As such, the dbcat combined effect is that the significant 1.3 mm detection confirms the marginal detection at 3 mm, otherwise these sources could have been missed at 3 mm mbcat. As shown in the two left panels in Fig. 4, the dbcat sources have detections with fluxes down to 0.01 mJy at 3 mm, about one order of magnitude fainter compared to mbcat. The second case is that when a source evolves into an UCHii region, the ionized gas produces strong free-free emission which flattens the spectral energy distribution at the millimeter wavelength. In this case, the 3 mm emission can be strongly enhanced, while the dust emission at 1.3 mm can be even weaker due to dust depletion by sputtering or sublimation and therefore becomes even more undetectable (Hoare et al. 1991; Inoue 2002; Draine 2011). For example, in Sgr C-Main, there are two significant detections SgrC-3mm#12 (Sig140) and SgrC-3mm#17 (Sig101), and are either marginally detected as SgrC-1mm#211 (Sig5) or not detected at all. But in the dbcat, both of them are detected as SgrC-db#25 (combined Sig83) and SgrC-db#30 (combined Sig65), respectively.
To summarize, mbcats are optimized for monochromatic completeness, while dbcat has improved robustness by dual-band cross check.
@@ 5 Intrinsic source size
Recent thousand-of-au interferometric observations have detected a large number of beam-like sources in distant and massive protoclusters (e.g., Liu et al. 2019; Motte et al. 2022; Xu et al. 2024b). After deconvolution, their intrinsic sizes are generally smaller than the beam. Are they objects with a uniform surface density across 1000 au or compact disks surrounded by less dense envelopes? Our DUET source catalog provides a large number of beam-like sources in a uniform spatial resolution. In Sect. 5.1, we estimate the intrinsic source size and orientation after deconvolution, demonstrating that most sources are unresolved. Furthermore, in Sect. 5.2, we use flux measurements to argue that the sources cannot be, or at least are not solely, point-like. Instead, they have finite source sizes but are still smaller than the beam.
@@ 5.1 Source deconvolution
The measured source size is the result of the intrinsic source size convolved with the beam size. Using the analytical method introduced in Appendix C, we can directly obtain the intrinsic major and minor axes as well as the position angle of the sources. We take dbcat as an example. As shown in the left panel of Fig. 5, most measured source sizes are larger than the beam size for both bands. After beam deconvolution, 77% sources have intrinsic sizes smaller than the beam size, which are referred to as marginally resolved sources. The distributions of deconvolved sizes of the 3 mm and 1.3 mm sources are more consistent between the two bands than the distributions of measured sizes, which indicates that the deconvolution successfully recovers the intrinsic source size. In the right panel, the observed source position angles show a clear concentration at the beam position angle, while the deconvolved position angles have a more scattered distribution. We interpret this as the convolution effects: the intrinsic source sizes should be smaller than or comparable to the beam size, so the beam-convolved position angles have a consistent beam-like direction. Once deconvolved from the beam, the tendency diminishes naturally.
5.2 Flux measurements
@@
In the mbcats, all the observed source sizes at 1.3 mm are larger than the beam size, whereas only one at 3 mm is point-like. In the dbcat, more 3 mm sources that are faint and blended can be identified with the help of 1.3 mm, including seven point-like sources at the 3 mm low flux end. Comparing between the top left and bottom left panels in Fig. 4, we find that the measured source fluxes in dbcat, especially at 3 mm, show a higher flux dynamic range than in mbcat. The limited flux dynamic range of the 3 mm sources is a direct result of the image sensitivity. However, when combined with 1.3 mm images, the detection sensitivity for sources at 3 mm can be improved from 0.1 mJy (10) down to 0.01 mJy (1). As discussed in Sect. 4.2.3, these faint 3 mm sources would probably be missed in monochromatic detection due to their low SNRs, which highlights the advantage of the dbcat compared to the mbcats.
@@ 6 Spectral indices
The spectral (intensity or brightness) index is defined as
where is the spectral intensity or brightness and is the frequency. Here, the subscripts (1,2) represent 1.3 mm and 3 mm observations, respectively.
Without considering dust scattering opacity, solving the equation of radiation transfer yields the following expression of spectral profile of dust thermal emission:
Here is the monochromatic blackbody emission at the temperature , which is written as,
@@
where is the monochromatic dust optical depth and can be expressed by the product of the monochromatic dust absorption opacity and the dust surface density . In the optically thin case, Eq. (16) gives assuming a dust emissivity spectral index 444The value is fitted from dust-ice mental models as described in Ossenkopf & Henning (1994). However, ALMA observations of high-mass protostellar cores generally show surface densities g cm-2 enclosed in the au scale (e.g. Motte et al. 2018a; Traficante et al. 2023; Xu et al. 2024b, c). We adopt the opacity model of protostellar dust grain with ice mantle in the – cm-3 density regime (Ossenkopf & Henning 1994) and obtain cm2/g and cm2/g. Assuming a gas-to-dust ratio , the dust optical depths are then obtained as and . As increases, converges toward , and the spectral index goes to 2.0 according to Eq. (17).
6.1 Low spectral indices or low brightnesses?

### v2

@@ 4.1 Source extraction strategy
The continuum images contain multiscale structures ranging from beam scales (02–03) to maximum recoverable scales (7″–14″), and exhibit a high intensity dynamic range up to 1000 (c.f. Col. 7 in Table 2). The getsf (v231026) extraction algorithm (Men’shchikov 2021) was adopted to identify and characterize sources, as it can identify sources from multi-scale and also from multiband data. We adopted the getsf definition of sources, which are roundish and significantly stronger than local surrounding fluctuations (of background and noise). The getsf algorithm is recommended for interferometric data, as it effectively handles various backgrounds from filamentary structures to negative bowls originated from imperfect uv-sampling and can separate blended sources in crowded fields like protoclusters (Xu et al. 2023; Cheng et al. 2024).
Our getsf procedure was designed to take full advantages of the dual-band observations. Details of the settings and the main workflow can be found in Appendix B. Here we only provide a brief overview. The inputs are the 1.3 mm and 3 mm continuum images at the original resolutions to avoid potential loss of detection due to the smoothing process. The images are without primary beam corrections for a more uniform noise distribution. The extraction mask includes regions where the primary beam response is greater than 0.3 (pb0.3), to exclude ambiguous detections on the edges. The minimum source size is set as the original resolution, and the maximum source size was set as five times the 1.3 mm beam size, 1.2 (approximately 0.05 pc or 10,000 au at the CMZ distance). The choice of maximum source size can not only effectively exclude those extended sources that are not core-like structures, but also preserve the completeness of the core-like source catalog. The main getsf procedure was divided into three rounds: preparing flattened components across the two bands; performing single-band extractions for monochromatic catalogs; and rerunning a dual-band detection to obtain a combined catalog.
@@ 4.2 Catalog description
Our procedure (see Sect. 4.1) produced two types of source catalogs: mono-band and dual-band. The mono-band source catalogs (mbcat) provide detections and measurements at individual bands, while the dual-band source catalogs (dbcat) combine sources333getsf accumulates the clean images (background removed) over dual bands by a natural weight, as is described by Eq. (29) in Men’shchikov (2021). decomposed from the two bands for detection and measurement. In the following, we introduce the mono-band catalogs in Sect. 4.2.1 and the dual-band catalogs in Sect.4.2.2, and discuss their caveats and complementarity in Sect. 4.2.3. The source catalogs are publicly available at 10.5281/zenodo.15063109.
@@ 4.2.1 Mono-band source catalogs
The 1.3 mm and 3 mm mbcats are presented in Tables 3 and 4, respectively. The significance value (Sig) is equivalent to the signal-to-noise ratio (S/N) of a source on the scale where it is best detected by getsf. The sources are sorted by the Sig value in the descending order, as indicated by the ID in the table. Throughout the DUET series, we use the nomenclature of “{cloud name}-{band}{#ID}”, where “cloud name” refers to either of cloude, SgrC, and 20kms. For example, SgrC-1mm#1 refers to the source with ID of 1 in the 1.3 mm mbcat in the Sgr C cloud, and 20kms-3mm#2 refers to the source with ID of 2 in the 3 mm mbcat in the 20 km s-1 cloud. Following the convention from Motte et al. (2018b), Sig greater than 7 corresponds to a robust detection. There are 437 and 278 robust detections at 1.3 mm and 3 mm in mbcats, corresponding to about 77% and 84% of the monochromatic detections, respectively. To obtain the intrinsic source sizes, we deconvolved the synthesized beams from the observed sizes by the analytical solution introduced in Appendix C. The deconvolved FWHM sizes were converted into physical scales at the distance of 8.277 kpc. The peak intensity and total flux were both primary-beam-corrected; the total flux of source is aperture-corrected by the radio beam shape rather than directly summarizing the flux per pixel that belongs to the source.
| ID | R.A. | Dec. | Sig | PA | Size | |||
| (J2000) | (J2000) | () | (deg) | (au) | (mJy beam-1) | (mJy) | ||
| 20 km s-1 | ||||||||
| Sgr C | ||||||||
| cloud e | ||||||||
| ID | R.A. | Dec. | Sig | PA | Size | |||
| (J2000) | (J2000) | () | (deg) | (au) | (mJy beam-1) | (mJy) | ||
| 20 km s-1 | ||||||||
| Sgr C | ||||||||
| cloud e | ||||||||
We did not perform any further filtering criteria based on their intensity profiles, to maintain mbcat’s completeness for future studies at individual bands. For example, some diffuse and extended sources can be analogous to nearby prestellar cores with sizes of 10,000 au (see Table 4 of Xu et al. 2024a, and references therein). The intensity profile of these prestellar cores should be much shallower than typically observed in protostellar cores because of a shallower density profile and a positive temperature gradient (colder in the inner part). Due to the missing flux effect, prestellar cores may appear even fainter than expected.
@@ 4.2.2 Dual-band source catalogs
The dbcat catalogs were obtained in the third round of getsf procedure (see Appendix B), where sources were identified and measured with the combination of both 1.3 mm and 3 mm continuum emissions. Since the FoVs of the two bands only partially overlap, the initial dbcats include sources outside the pb=0.3 mask of one of the bands. We removed such sources to obtain the cleanest source sample within the overlapping FoVs. Finally, a total of 450 sources (including 62 in cloud e, 182 in Sgr C, and 206 in the 20 km s-1 cloud) were included in the dbcat; they are listed in Table 6. The complete version of dbcat is shared in 10.5281/zenodo.15068333. Throughout the paper, we use the nomenclature of “{cloud name}-{db}#ID”. For example, SgrC-db#1 refers to the source with ID of 1 in the dbcat in the Sgr C cloud. As an example, the spatial distribution of the identified sources in the Sgr C cloud is shown in Fig. 3. The complete version for all the three clouds are shared both in an interactive mode (https://xfengwei.github.io/magnifier/index.html) and in 10.5281/zenodo.15068333.
The dbcat sources have undergone a similar filtering criteria as the mbcat ones, but the goodness and significance are combined from both 1.3 mm and 3 mm. Therefore, some dbcat sources could have a small Sig value in one band because of a significant detection in the other band. For example, SgrC-db#30 is marginally detected at 1.3 mm but filtered by Sig and goodness criteria, while it has S/N65 at 3 mm. Another example is SgrC-db#106, which has a S/N17 detection at 1.3 mm but marginal detection at 3 mm. In the complete version of Table 6, the monochromatic Sig values at 1.3 mm and 3 mm are also given.
The dbcat can also resolve crowded sources thanks to the higher angular resolution at 1.3 mm. We refer to the cloud e-Main as an example. We identified cloude-1mm#1, 15, 21, 23, 28, 29, 55, and 61 from a clustered region at 1.3 mm, but only cloude-3mm#1, 16, and 22 at 3 mm due to its coarser resolution. However, all eight sources were successfully separated and identified in the dbcat.
Source sizes and fluxes at each band were measured within individual monochromatic footprints, so the sizes can differ between the two bands. Two data columns for 1.3 mm and 3 mm, respectively, were include in the catalog, each of which contains the same parameters as the mbcats. The intrinsic source sizes of both bands were obtained by deconvolving the synthesized beams from the observed sizes (see Appendix C). In Table 6 listed are the major and minor FWHM sizes, position angle, intrinsic source size, peak intensity, and integrated flux over the source.
@@ 4.2.3 Completeness and robustness of the catalogs
As is shown by solid orange and yellow lines in Fig. 1, the FoVs at 1.3 mm and 3 mm only partially overlap. The mbcats contain all the detections within the FoVs at the individual bands, while the dbcat only includes sources appearing in the overlapped FoVs of the two bands. In this sense, the mbcat ensures the completeness of the sources within the FoV at each individual band.
Meanwhile, dbcat considers both bands simultaneously for detection, raising the robustness of source detection. For example, the mass sensitivity of optically thin dust at 20 K at 1.3 mm is about four times better than 3 mm, so a faint dusty source can have a robust detection at 1.3 mm but less than at 3 mm. In this case, the 1.3 mm mbcat will include this source, but the dbcat will not. Therefore, the 1.3 mm mbcat catalog is more complete than the dbcat at the low-flux end. In turn, sources in the dbcat are more robust, because they are verified in both bands. This is particularly important for interferometric images, where there can be sidelobes or artifacts from cleaning. Compared to the mbcat, the dbcat combines images to filter out false detections, resulting in a smaller but more robust source catalog.
However, it should be noted that if a source is marginally detected and filtered out in one band but has a robust detection in another, the dbcat will contain it. It usually happens in two cases. The first case is also for optically thin dust core mentioned above, but the difference happens when a 1.3 mm detection reaches as much as 20 but 3 mm is still below 5. As such, the combined dbcat effect is that the significant 1.3 mm detection confirms the marginal detection at 3 mm, otherwise these sources could have been missed at 3 mm mbcat. As is shown in the two left panels in Fig. 4, the dbcat sources have detections with fluxes down to 0.01 mJy at 3 mm, about one order of magnitude fainter compared to mbcat. The second case is that when a source evolves into an UCHii region, the ionized gas produces strong free-free emission which flattens the spectral energy distribution at the millimeter wavelength. In this case, the 3 mm emission can be strongly enhanced, while the dust emission at 1.3 mm can be even weaker due to dust depletion by sputtering or sublimation and therefore becomes even more undetectable (Hoare et al. 1991; Inoue 2002; Draine 2011). For example, in Sgr C-Main, there are two significant detections SgrC-3mm#12 (Sig140) and SgrC-3mm#17 (Sig101), and are either marginally detected as SgrC-1mm#211 (Sig5) or not detected at all. But in the dbcat, both of them are detected as SgrC-db#25 (combined Sig83) and SgrC-db#30 (combined Sig65), respectively. To summarize, mbcats are optimized for monochromatic completeness, while dbcat has improved robustness by dual-band cross check.
@@ 5 Intrinsic source size
Recent interferometric observations on the order of thousands of astronomical units have detected a large number of beam-like sources in distant and massive protoclusters (e.g., Liu et al. 2019; Motte et al. 2022; Xu et al. 2024b). After deconvolution, their intrinsic sizes are generally smaller than the beam. A question arises whether they are objects with a uniform surface density across 1000 au or compact disks surrounded by less dense envelopes. Our DUET source catalog provides a large number of beam-like sources in a uniform spatial resolution. In Sect. 5.1, we estimate the intrinsic source size and orientation after deconvolution, demonstrating that these sources are mostly unresolved. Furthermore, in Sect. 5.2, we use flux measurements to argue that the sources cannot be, or at least are not solely, point-like. Instead, they have finite source sizes but are still smaller than the beam.
@@ 5.1 Source deconvolution
The measured source size is the result of the intrinsic source size convolved with the beam size. Using the analytical method introduced in Appendix C, we can directly obtain the intrinsic major and minor axes as well as the position angle of the sources. We take dbcat as an example. As is shown in the left panel of Fig. 5, most measured source sizes are larger than the beam size for both bands. After beam deconvolution, 77% sources have intrinsic sizes smaller than the beam size; these are referred to as marginally resolved sources. The distributions of deconvolved sizes of the 3 mm and 1.3 mm sources are more consistent between the two bands than the distributions of measured sizes, which indicates that the deconvolution successfully recovers the intrinsic source size. In the right panel, the observed source position angles show a clear concentration at the beam position angle, while the deconvolved position angles have a more scattered distribution. We interpret this as the convolution effects: the intrinsic source sizes should be smaller than or comparable to the beam size, so the beam-convolved position angles have a consistent beam-like direction. Once deconvolved from the beam, the tendency diminishes naturally.
5.2 Flux measurement
@@
In the mbcats, all the observed source sizes at 1.3 mm are larger than the beam size, whereas only one at 3 mm is point-like. In the dbcat, more 3 mm sources that are faint and blended can be identified with the help of 1.3 mm, including seven point-like sources at the 3 mm low flux end. Comparing between the top left and bottom left panels in Fig. 4, we find that the measured source fluxes in dbcat, especially at 3 mm, show a higher flux dynamic range than in mbcat. The limited flux dynamic range of the 3 mm sources is a direct result of the image sensitivity. However, when combined with 1.3 mm images, the detection sensitivity for sources at 3 mm can be improved from 0.1 mJy (10) down to 0.01 mJy (1). As is discussed in Sect. 4.2.3, these faint 3 mm sources would probably be missed in monochromatic detection due to their low S/Ns, which highlights the advantage of the dbcat compared to the mbcats.
@@ 6 Spectral indices
The spectral index is defined as
where is the monochromatic intensity or brightness at the frequency . In this work, we are discussing the inter-band spectral index where its subscripts () represent 1.3 mm and 3 mm, respectively. By solving the equation of radiation transfer without dust scattering yields
where is the monochromatic blackbody emission at the temperature ,
and is the monochromatic dust optical depth and can be expressed by the product of the monochromatic dust absorption opacity, , and the dust surface density, .
@@
We discuss in different regime. In the optically thin case, Eq. (16) gives assuming a dust emissivity spectral index 666The value is fitted from dust-ice mental models as described in Ossenkopf & Henning (1994). However, recent ALMA observations of high-mass protostellar cores have shown surface densities g cm-2 within au scale (e.g., Motte et al. 2018a; Traficante et al. 2023; Xu et al. 2024b, c), which makes dust optically thick. We adopt the opacity model of protostellar dust grain with ice mantle (see Fig. 12) in the – cm-3 density regime (Ossenkopf & Henning 1994) and obtain cm2/g and cm2/g. Assuming a gas-to-dust ratio of , the dust optical depths are then obtained as and . As increases, converges toward , and the spectral index goes to 2.0 according to Eq. (17).
6.1 Low spectral indices versus low brightnesses